# Fuzzy Information Evolution with Three-Way Decision in Social Network Group Decision-Making

### v1

@@ IV Example
In recent years, with the development of multi-UAV cooperative technology, UAV swarm intelligence systems have demonstrated a high level of autonomy and efficient information-sharing capabilities across various fields. It is worth noting that this study is inspired by the research presented in [34]. As shown in Fig. 1, a group of $n$ UAVs is cooperatively circumnavigating a group of $m$ unknown targets on a horizontal plane. Taking post-earthquake rescue in mountainous areas as an example, multiple UAVs are often deployed to search for survivors across large regions. In this setting, the model based on collaborative decision-making and opinion dynamics is essential. Each UAV acts as an intelligent agent, and shares the linguistic information with the system. As these agents communicate, they update the evaluation information of the situation.
@@
To evaluate the proposed linguistic 3WD opinion dynamics model, we conduct simulations with a fixed number of agents and predefined model parameters. Specifically, the number of agents was set to $N=20$. The linguistic term set consisted of $12\Phi+$ 1 elements, with $\Phi=3$, and the non-linear mapping parameter was set to $a=2$. The 3WD is controlled by two thresholds: the strict acceptance threshold $\alpha=0.3$, and the strict rejection threshold $\beta=0.6$. Agents falling in the uncertain region $(\alpha,\beta)$ were accepted with a probability that decays exponentially with the distance from $\alpha$, governed by a decay factor $\lambda=10$. The opinion inertia parameter was set to $\omega=0$, meaning that agents fully adopt the weighted average of their accepted neighbors’ opinions without retaining their previous value. For the dynamic network update process, we adopted an additive threshold $\delta_{\text{add}}=0.15$ and a cut-off threshold $\delta_{\text{cut}}=0.45$. Links were added or removed with probabilities $p_{\text{add}}=0.5$ and $p_{\text{cut}}=0.5$, respectively. The simulation ran for at most $T_{\max}=10$ time steps, or until convergence was reached, defined by a maximum opinion change less than a tolerance of $\varepsilon=10^{-3}$.
@@
Fig. 2 shows how the opinions of all agents change over time. We can see that several agents adjusted their opinions clearly in the early steps, which means that there is active communication and belief change at the beginning. Most of these changes happened within the first four iterations. After that, the opinions became more stable. By the fifth iteration, most agents had nearly stopped changing. At the seventh iteration, all opinions stayed the same, which means the system reached a stable state and no more updates are needed. The final linguistic term for each agent is shown in Fig. 3. Fig. 4 presents the evolution of the social network and agent opinions from iteration 0 to 7. In each subplot, nodes represent agents and edges represent social ties. Node colors indicate the linguistic term each agent holds at that time step. It can be seen that agents with similar opinions gradually cluster together as the simulation progresses. From iteration 5 onwards, the network evolves into two densely connected subgroups, each associated with an opinion. Fig. 5 shows how the network structure changes over the iterations. The average degree increases quickly from about 1.9 to nearly 12, which means that agents are building more links with each other as iteration goes on. This suggests that agents are becoming more similar in their opinions, so new connections are more likely under the network update rule. Meanwhile, the number of isolated nodes drops from 12 to 0 in the first two steps and stays at zero after that. This means that all agents are connected to the network early on, and the structure remains fully connected during the process. These results show that the network becomes more connected over time.
@@ V Comparison, Robustness Analysis, and Complexity Analysis
V-A Impact of Broad Acceptance Thresholds on Consensus Formation
@@
In 3WD framework where $[\alpha,\beta]$ is the hesitation region, the choice of thresholds directly influences how likely agents are to exchange information. In Fig. 6, $\alpha$ is set to 0.2 and $\beta$ to 0.4, which defines a narrow acceptance range. Under this setting, only a small number of opinion differences fall within the acceptance or hesitation zones, while most are rejected directly. As a result, agents are more likely to avoid opinion exchange, leading to more isolated updates and a wider spread in final linguistic term indices. This produces a more diverse and polarized results. In contrast, Fig. 3 uses $\alpha=0.3$ and $\beta=0.6$, expanding both the acceptance and hesitation regions. More agent pairs are allowed to interact, either certainly or probabilistically, which leads to smoother opinion integration. Therefore, smaller $\alpha$ and $\beta$ reduce communication opportunities and promote diversity, while larger thresholds encourage interaction and support consensus formation.
In Fig. 7, the thresholds are set to $\alpha=0.5$ and $\beta=0.9$, which significantly enlarges both the hesitation and acceptance regions compared to the previous settings. Under this configuration, only opinion differences greater than 0.9 are directly rejected, so most agent pairs have a chance to exchange information. As shown in the figure, the majority of agents reach a high and consistent linguistic term index $h_{3}$, while only a few remain at the lowest level $h_{0}$. This result indicates that a wider acceptance rule allows more interaction among agents, helping them to integrate their opinions. Compared to smaller values of $\alpha$ and $\beta$, this setup is more likely to support global agreement.
@@
V-B Impact of Network Scale on Opinion Evolution and Consensus Formation
Figs. 4, 8, and 9 respectively illustrate the evolution process of the system when the number of agents is 20, 30, and 50, under the proposed model with 3WD mechanism and dynamic social network reconstruction rules. As the number of agents increases, the initially sparse social network gradually evolves into several closely connected subgroups. When $N=20$, the connections between agents are relatively concentrated, and two consensus groups are eventually formed. When $N=30$, the trend of network division becomes stronger, and some agents with extreme opinions form isolated small groups disconnected from the main network. In the large-scale system with $N=50$, although the initial connections are very sparse, the main group gradually absorbs most agents through the self-adjusting structure of the network, and eventually forms a unified group with closer connections and more consistent opinions. This phenomenon fully demonstrates the effectiveness and advantages of the proposed algorithm. As the scale of the social network increases, differences between agents grow, and the uncertainty of connections rises. Traditional static network models are likely to face problems such as opinion polarization, information isolation, and failure to reach consensus. In contrast, our model introduces a dynamic linking mechanism based on opinion distance and a “stay undecided” option from the 3WD theory, which flexibly handles agents near the decision boundary and significantly improves the system’s ability to reach consensus under complex conditions. As shown in the figures, this mechanism not only helps to reduce the trend of division caused by sparse networks and opinion differences but also enhances information flow and coordinated evolution within the group.
@@ V-B Impact of Network Scale on Opinion Evolution and Consensus Formation
V-C Impact of Three-Way Decision Mechanism on Opinion Dynamics
To evaluate the effect of the proposed 3WD mechanism on the evolution of group opinions, we conducted a comparative study involving four scenarios. Fig. 10 shows a comparison of convergence speed in four different scenarios. Each subplot represents a specific setting, where the horizontal axis is the number of iterations and the vertical axis shows the maximum change in opinions per step, denoted as $\delta_{\text{max}}$. This value is used as the measurement of convergence speed in all four experiments, and it is calculated by the following formula:
@@ V-C Impact of Three-Way Decision Mechanism on Opinion Dynamics
Here, $\theta_{i}(t)$ represents the opinion of agent $i$ at $t$. This metric reflects the largest individual change at each step and helps evaluate how fast or smooth the opinion evolution process is. In the experiments, Scenario 1 and Scenario 3 represent the cases with 40 and 60 agents respectively, where 3WD mechanism is applied using parameters $\alpha=0.3$ and $\beta=0.6$. Scenario 2 and Scenario 4 use the same number of agents but do not include the 3WD mechanism, which is represented by $\alpha=\beta=0.6$, meaning no hesitation zone is added. From the figure, we can see that in Scenario 2 and Scenario 4 (without the mechanism), $\delta_{\text{max}}$ drops quickly, and convergence is reached within 4 and 6 iterations respectively. The process is fast and smooth. In contrast, Scenario 1 and Scenario 3 (with the mechanism) show a slower convergence process, with a more gradual decrease in $\delta_{\text{max}}$ and some fluctuations, especially in Scenario 3. Overall, the 3WD mechanism slows down the speed of opinion convergence, which results in a slower but more stable evolution. This effect becomes more visible when the number of agents increases. These results suggest that the 3WD mechanism achieves a balance between convergence speed and system stability.
@@
V-D Impact of $\delta_{\text{add}}$ on Opinion Dynamics
@@
V-E Impact of $\delta_{\text{cut}}$ on Opinion Dynamics
@@
V-F Complexity Analysis
V-F1 Time Complexity
@@
In each iteration, the proposed model checks all possible neighbor pairs $(i,j)$ to compute opinion distances and determine acceptance. In the worst case, each agent needs to compare with $N-1$ others, resulting in a complexity of $\mathcal{O}(N^{2})$ for this step. The dynamic rewiring step also involves checking all agent pairs and may update the network structure, which again takes $\mathcal{O}(N^{2})$ time. The convergence check requires scanning all agents, which takes $\mathcal{O}(N)$ time. Therefore, each iteration has a total time complexity of $\mathcal{O}(N^{2})$, and with a maximum of $T_{\max}$ iterations, the overall time complexity is $\mathcal{O}(T_{\max}\cdot N^{2})$.
V-F2 Space Complexity
@@
The main memory usage comes from storing the network and agent states. The adjacency matrix takes $\mathcal{O}(N^{2})$ space. Each agent’s numerical opinion $\theta_{i}(t)$ and linguistic opinion $x_{i}(t)$ together take $\mathcal{O}(N)$ space. In addition, the neighbor sets may each contain up to $N$ elements, which leads to a worst-case memory usage of $\mathcal{O}(N^{2})$. Therefore, the total space complexity is $\mathcal{O}(N^{2})$.
@@
In summary, this algorithm models the information evolution with a reasonable level of computational complexity. Although the worst-case time and space complexity are both $\mathcal{O}(N^{2})$, real-world social networks are usually sparse and locally connected, so the actual runtime is often much lower than the theoretical upper bound. Moreover, the model has a clear structure and is easy to implement in parallel.
V-G Potential Applications
V-G1 Brain-Computer Interface (BCI) systems

### v2

@@ IV Example
In recent years, with the development of multi-UAV cooperative technology, UAV swarm intelligence systems have demonstrated a high level of autonomy and efficient information-sharing capabilities across various fields. It is worth noting that this study is inspired by the research presented in [34]. As shown in Fig. 1, a group of $n$ UAVs is cooperatively circumnavigating a group of $m$ unknown targets on a horizontal plane. In this paper, we take post-earthquake rescue in mountainous areas as an example. Multiple UAVs are often deployed to search for survivors across large regions. In this setting, the model based on collaborative decision-making and opinion dynamics is essential. Each UAV acts as an intelligent agent and shares the linguistic information with the system. As these agents communicate, they update the evaluation information of the situation.
@@
To evaluate the proposed linguistic 3WD opinion dynamics model, we conduct simulations with a fixed number of agents and predefined model parameters. Specifically, the number of agents was set to $N=20$. The linguistic term set consisted of $2\Phi+1$ elements, with $\Phi=3$, and the non-linear mapping parameter was set to $a=2$. The 3WD is controlled by two thresholds: the strict acceptance threshold $\alpha=0.3$, and the strict rejection threshold $\beta=0.6$. Agents falling in the uncertain region $(\alpha,\beta)$ were accepted with a probability that decays exponentially with the distance from $\alpha$, governed by a decay factor $\lambda=10$. The opinion inertia parameter was set to $\omega=0$, meaning that agents fully adopt the weighted average of their accepted neighbors’ opinions without retaining their previous value. For the dynamic network update process, we adopted an additive threshold $\delta_{\text{add}}=0.15$ and a cut-off threshold $\delta_{\text{cut}}=0.45$. Links were added or removed with probabilities $p_{\text{add}}=0.5$ and $p_{\text{cut}}=0.5$, respectively. The simulation ran for at most $T_{\max}=10$ time steps, or until convergence was reached, defined by a maximum opinion change less than a tolerance of $\varepsilon=10^{-3}$.
@@
Fig. 2 shows how the opinions of all agents change over time. We can see that several agents adjusted their opinions clearly in the early stages, which indicates active communication and belief change at the outset. Most of these changes happened within the first four iterations. After that, the opinions became more stable. By the fifth iteration, most agents had nearly stopped changing. At the seventh iteration, all opinions stayed the same, which means the system reached a stable state and no more updates are needed. The final linguistic term for each agent is shown in Fig. 3. Fig. 4 presents the evolution of the social network and agent opinions from iteration 0 to 7. In each subplot, nodes represent agents and edges represent social ties. Node colors indicate the linguistic term each agent holds at that time step. It can be seen that agents with similar opinions gradually cluster together as the simulation progresses. From iteration 5 onwards, the network evolves into two densely connected subgroups, each associated with an opinion. Fig. 5 shows how the network structure changes over the iterations. The average degree increases quickly from about 1.9 to nearly 12, which means that agents are building more links with each other as the iteration goes on. This suggests that agents are becoming more similar in their opinions, so new connections are more likely under the network update rule. Meanwhile, the number of isolated nodes drops from 1 to 0 in the first two steps and stays at zero after that. This means that all agents are connected to the network early on, and the structure remains fully connected during the process. These results show that the network becomes more connected over time.
To quantitatively evaluate the consensus degree of the system, this paper adopts three complementary indicators. First, the opinion variance $\mathrm{Var}(t)=\frac{1}{n}\sum_{i=1}^{n}\bigl(x_{i}(t)-\bar{x}(t)\bigr)^{2}$ is used to reflect the overall dispersion of the group, where a smaller variance indicates stronger convergence. Second, the opinion range $R(t)=\max_{i}x_{i}(t)-\min_{i}x_{i}(t)$ is applied to measure the extreme difference within the group, and $R(t)=0$ means that all agents have reached complete consensus. Finally, we introduce the normalized consensus index $C_{\mathrm{AAD}}(t)=1-\frac{1}{n}\sum_{i=1}^{n}\frac{|x_{i}(t)-\bar{x}(t)|}{d_{\max}}$ based on the average absolute deviation, and $d_{\max}$ is the maximum possible deviation in the opinion space. This index is normalized to the interval $[0,1]$, and values closer to 1 indicate a higher level of consensus. By combining these three measures, the consensus level of the system can be comprehensively evaluated from the perspectives of overall convergence, extreme differences, and normalized description. In this case, we can obtain $\mathrm{Var}(7)=0.0469$, $R(7)=0.5$, and $C_{\mathrm{AAD}}(7)=0.625$. These values show that the opinions still have some dispersion (since the variance 0.0469 is not close to 0), and the largest difference between individuals is 0.5, which means there is a clear gap. At the same time, the normalized consensus index is 0.625, which indicates that the system has not reached full agreement but rather a state of partial consensus.
@@ V Comparison, Robustness Analysis, and Complexity Analysis
V-A Impact of Parameters and Mechanisms on Proposed Model
V-A1 Impact of Broad Acceptance Thresholds on Consensus Formation
@@
In the 3WD framework, where $\alpha,\beta]$ is the hesitation region, the choice of thresholds directly influences how likely agents are to exchange information. In Fig. 6, $\alpha$ is set to 0.2 and $\beta$ to 0.4, which defines a narrow acceptance range. Under this setting, only a small number of opinion differences fall within the acceptance or hesitation zones, while most are rejected directly. As a result, agents are more likely to avoid opinion exchange, leading to more isolated updates and a wider spread in final linguistic term indices. This produces more diverse and polarized results. In contrast, Fig. 3 uses $\alpha=0.3$ and $\beta=0.6$, expanding both the acceptance and hesitation regions. More agent pairs are allowed to interact, either certainly or probabilistically, which leads to smoother opinion integration. Therefore, smaller $\alpha$ and $\beta$ reduce communication opportunities and promote diversity, while larger thresholds encourage interaction and support consensus formation.
In Fig. 7, the thresholds are set to $\alpha=0.5$ and $\beta=0.9$, which significantly enlarges both the hesitation and acceptance regions compared to the previous settings. Under this configuration, only opinion differences greater than 0.9 are directly rejected, so most agent pairs have a chance to exchange information. As shown in the figure, the majority of agents reach a high and consistent linguistic term index $h_{3}$, while only a few remain at the lowest level $h_{0}$. This result suggests that a wider acceptance rule enables more interaction among agents, facilitating their integration of opinions. Compared to smaller values of $\alpha$ and $\beta$, this setup is more likely to support global agreement.
@@
V-A2 Impact of Network Scale on Opinion Evolution and Consensus Formation
Figs. 4, 8, and 9 respectively illustrate the evolution process of the system when the number of agents is 20, 30, and 50, under the proposed model with 3WD mechanism and dynamic social network reconstruction rules. As the number of agents increases, the initially sparse social network gradually evolves into several closely connected subgroups. When $N=20$, the connections between agents are relatively concentrated, and two consensus groups are eventually formed. When $N=30$, the trend of network division becomes stronger, and some agents with extreme opinions form isolated small groups disconnected from the leading network. In the large-scale system with $N=50$, although the initial connections are very sparse, the leading group gradually absorbs most agents through the self-adjusting structure of the network. Eventually, it forms a unified group with closer connections and more consistent opinions. This phenomenon fully demonstrates the effectiveness and advantages of the proposed algorithm. As the scale of the social network increases, differences between agents grow, and the uncertainty of connections rises. Traditional static network models are likely to face problems such as opinion polarization, information isolation, and failure to reach consensus. In contrast, our model introduces a dynamic linking mechanism based on opinion distance and a “stay undecided” option from the 3WD theory, which flexibly handles agents near the decision boundary and significantly improves the system’s ability to reach consensus under complex conditions. As shown in the figures, this mechanism not only helps to reduce the trend of division caused by sparse networks and opinion differences but also enhances information flow and coordinated evolution within the group.
@@ V-A2 Impact of Network Scale on Opinion Evolution and Consensus Formation
V-A3 Impact of Three-Way Decision Mechanism on Opinion Dynamics
To evaluate the effect of the proposed 3WD mechanism on the evolution of group opinions, we conducted a comparative study involving four scenarios. Fig. 10 shows a comparison of convergence speed in four different scenarios. Each subplot represents a specific setting, where the horizontal axis is the number of iterations and the vertical axis shows the maximum change in opinions per step, denoted as $\delta_{\text{max}}$. This value is used as the measurement of convergence speed in all four experiments, and the following formula calculates it:
@@ V-A3 Impact of Three-Way Decision Mechanism on Opinion Dynamics
Here, $\theta_{i}(t)$ represents the opinion of agent $i$ at $t$. This metric reflects the largest individual change at each step and helps evaluate how fast or smooth the opinion evolution process is. In the experiments, Scenarios 1 and 3 represent cases with 40 and 60 agents, respectively, where the 3WD mechanism is applied using parameters $\alpha=0.3$ and $\beta=0.6$. Scenarios 2 and 4 use the same number of agents but do not include the 3WD mechanism, which is represented by $\alpha=\beta=0.6$, meaning no hesitation zone is added. From the figure, we can see that in Scenarios 2 and 4 (without the mechanism), $\delta_{\text{max}}$ drops quickly, and convergence is reached within 4 and 6 iterations, respectively. The process is fast and smooth. In contrast, Scenarios 1 and 3 (with the mechanism) show a slower convergence process, with a more gradual decrease in $\delta_{\text{max}}$ and some fluctuations, especially in Scenario 3. Overall, the 3WD mechanism slows down the speed of opinion convergence, which results in a slower but more stable evolution. This effect becomes more visible when the number of agents increases. These results suggest that the 3WD mechanism achieves a balance between convergence speed and system stability.
@@
V-A4 Impact of $\delta_{\text{add}}$ on Opinion Dynamics
@@
V-A5 Impact of $\delta_{\text{cut}}$ on Opinion Dynamics
@@
V-B Complexity Analysis
V-B1 Time Complexity
@@
In each iteration, the proposed model checks all possible neighbor pairs $(i,j)$ to compute opinion distances and determine acceptance. In the worst case, each agent needs to compare with N1N-1 others, resulting in a complexity of $\mathcal{O}(N^{2})$ for this step. The dynamic rewiring step also involves checking all agent pairs and may update the network structure, which again takes $\mathcal{O}(N^{2})$ time. The convergence check requires scanning all agents, which takes $\mathcal{O}(N)$ time. Therefore, each iteration has a total time complexity of $\mathcal{O}(N^{2})$, and with a maximum of $T_{\max}$ iterations, the overall time complexity is $\mathcal{O}(T_{\max}\cdot N^{2})$.
V-B2 Space Complexity
@@
The primary source of memory usage is storing the network and agent states. The adjacency matrix takes $\mathcal{O}(N^{2})$ space. Each agent’s numerical opinion $\theta_{i}(t)$ and linguistic opinion $x_{i}(t)$ together take $\mathcal{O}(N)$ space. In addition, the neighbor sets may each contain up to $N$ elements, which leads to a worst-case memory usage of $\mathcal{O}(N^{2})$. Therefore, the total space complexity is $\mathcal{O}(N^{2})$.
@@
To test how computation time changes with the number of agents, we carried out experiments on the MATLAB 2020a platform, and the results are shown in Fig. 13. Between 10 and 20 agents, the computation time increases clearly, from about 3.5 seconds to nearly 7.8 seconds. When the number of agents rises from 20 to 25, the time drops slightly to about 6.7 seconds. After that, between 25 and 35 agents, the time increases again and reaches about 8.2 seconds at 35 agents. These results indicate that as the number of agents increases, the overall computation time grows, but the fluctuation is not significant and the growth is not strictly linear. In summary, this algorithm models the information evolution with a reasonable level of computational complexity. Although the worst-case time and space complexity are both $\mathcal{O}(N^{2})$, real-world social networks are usually sparse and locally connected, so the actual runtime is often much lower than the theoretical upper bound. Moreover, the model has a clear structure and is easily implementable in parallel.
Figure 13: Impact of Agent Number on Computation Time.
V-C Comparative Analysis with Mainstream Models
In response to the reviewer’s suggestion, we conducted additional comparative experiments with several representative opinion dynamics models. The aim was to evaluate the proposed model against widely used approaches and to verify whether the observed behaviors are consistent or significantly different. All simulations in this section use the same initial opinions as those described in Section Example.
V-C1 DeGroot Model
The DeGroot model is a classical linear consensus model in which each agent updates its opinion by taking a weighted average of all agents’ opinions, as introduced in Definition 3. Here, two variants of the DeGroot model are considered:
  • Uniform weights: In this case, each agent assigns the same weight to every other agent, regardless of the opinion distance. This represents the most basic form of the DeGroot update rule.
  • Distance-based weights: In this variant, the interaction weight between two agents depends on the distance between their opinions. The weight matrix is calculated using $\omega_{j}=\frac{e^{-|D_{ij}|}}{\sum_{k=1}^{20}e^{-|D_{ik}|}}$. This formulation ensures that agents with more similar opinions exert a stronger influence on each other.
Under the DeGroot model, we compare two methods: uniform weights and distance-based weights. The results are shown in Fig. 14. In the uniform weight case, the weight of each agent is the same, 1/20. Since each row of the weight matrix is identical, the system quickly converges after the first iteration to the arithmetic mean of all initial opinions. It remains unchanged, showing an averaging effect and one-step convergence. In the distance-based weight case, agents with higher opinion similarity have a stronger influence on each other. In the early stage, agents with extreme opinion values change faster, while those in the middle change more slowly. The system becomes stable after 2–3 iterations, and the final consensus value is the weighted average of all initial opinions, which is slightly different from the simple average in the uniform weight case. These results show that the DeGroot model is sensitive to the way the weights are calculated.
V-C2 Homogeneous HK Bounded Confidence Model
The homogeneous HK bounded confidence model assumes that all agents share the same confidence bound $\epsilon$, as introduced in Definition 4. At each iteration, an agent only considers the opinions of others whose opinion difference is less than $\epsilon$. To enable a thorough comparative analysis, we tested six different values of $\epsilon$, namely $0.35$, $0.30$, $0.25$, $0.20$, $0.15$, and $0.10$. As $\epsilon$ decreases, the communication threshold between agents becomes smaller, meaning that an agent interacts only with those whose opinions are very similar to its own. Fig. 15 illustrates the opinion evolution of all agents under six different confidence bound values in the homogeneous HK bounded confidence model. When $\epsilon=0.35$ [Fig. 15(a)], several agents are within each other’s confidence intervals, leading to active opinion exchange across multiple groups. This interaction results in partial convergence toward intermediate values after a few iterations. When $\epsilon=0.30$ [Fig. 15(b)], the interaction range slightly decreases. However, since most of the key interaction links remain, the evolution results are the same as in the $\epsilon=0.35$ case. For $\epsilon=0.25$ [Fig. 15(c)], the threshold is further reduced, causing only a subset of agents to interact. Opinion exchange occurs primarily within smaller local clusters, and the number of isolated agents increases. When $\epsilon=0.20$ [Fig. 15(d)], $\epsilon=0.15$ [Fig. 15(e)], and $\epsilon=0.10$ [Fig. 15(f)], as the threshold gradually decreases, the information exchange between agents becomes less frequent, making it difficult for them to reach consensus in the end. Through the analysis, it is evident that this model is susceptible to the threshold, where even slight differences can result in significant variations in the results.
V-C3 Heterogeneous HK Bounded Confidence Model
The heterogeneous HK bounded confidence model extends the homogeneous version by assigning different threshold values to agents. In the first case, the threshold $\epsilon_{i}$ of each agent is listed below:
[…]
In the second case, $\epsilon_{10}$ is changed from 0.3 to 0.2, while all other values remain unchanged. In the third case, $\epsilon_{17}$ is changed from 0.7 to 0.3, while all other values remain unchanged. The results are shown in Fig. 16. The final results are distinctly different. Case 1 converges to two distinct opinion clusters, Case 2 produces a different set of stable clusters, and Case 3 results in yet another opinion distribution. These results demonstrate that the model is also susceptible to individual parameter values, where even a small change can lead to significantly different results.
Figure 14: Opinion Evolution under the DeGroot Model with Uniform and Distance-Based Weights.
Figure 15: Opinion Evolution of All Agents under Different Confidence Bounds in the Homogeneous HK Model.
Figure 16: Opinion Evolution of All Agents under Different Confidence Bounds in the Heterogeneous HK Model.
V-D Potential Applications
V-D1 Brain-Computer Interface (BCI) systems